A vehicle travels 125.7 feet while its weels revolve 16 times.Find the diameter of the wheels to the nearest inch.: A vehicle travels 125.7 feet while its weels revolve 16 times.Find the diameter of the wheels to the nearest inch.
Answer by stanbon(19018)

A vehicle travels 125.7 feet while its weels revolve 16 times.Find the diameter of the wheels to the nearest inch.
-------------------
125.7/2(pi)r = 16
r = 125.7/(2(pi)16)
r = 1.2503... ft.
--------
diameter = 2r
==================
Cheers,
Stan H.
Question 141745: A vehicle travels 125.7 feet while its weels revolve 16 times.Find the diameter of the wheels to the nearest inch.: A vehicle travels 125.7 feet while its weels revolve 16 times.Find the diameter of the wheels to the nearest inch.
Answer by ankor@dixie-net.com(4538)

A vehicle travels 125.7 feet while its wheels revolve 16 times. Find the diameter of the wheels to the nearest inch.
:
Find the circumference:
125.7/16 = 7.85625 ft
:
Convert to inches:
7.85625 * 12 = 94.275 inches
:
We know C = 2*pi*r or just pi*d
pi*d = 94.275
d = 94.275/pi
d = 30 inches
:
Check solution:
distance = (pi*30*16)/12 = 125.6637 ~ 125.7 ft